• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 08:19:21 on September 12, 2009

    Entry number 289208

    keyword=Owl Shift Summary

    September 12, 2009
    SL: V. Sulkosky
    TO: X. Yan
    3rd: J. Roche

    Owl Shift Summary:
    +++++++++++++++++++++++++++++++++++++++++
    00:00 Begin shift with beam tuning. Compton photon rates at 500 Hz/uA after Moller.
    Right Q2 magnet is down due to power supply problems.
    01:05 Compton photon rate now between 100-200 Hz/uA.
    01:14 Beam profile looks good from Harp scans:
              A: 278 x 62 microns
              B: 310 x 156 microns
    01:16 Move to holey carbon target to check raster. Ask for raster on (11 x 8.5 MCC units).
    01:19 Target at holey carbon, asked for 5 uA to check raster size.
    01:30 Noticed BPMB positions were off. Asked MCC to set them to the nominal positions (-0.8,2.1).
    01:40 Beam positions look good now.
    01:43 Move target to loop 3 LH2 target.
    01:47 Target at LH2.
    01:54 S0 HV's turned off after raster check.
    02:10 Compton HV for GSO tripped, reset.
    02:15 Compton HV for GSO tripped again, reset.
    03:23 Changed fast FB BPM gain control from auto to manual.
    03:39 Change current to 80 uA during boiling study.
    03:55 Change current to 60 uA during boiling study.
    04:13 Change current to 40 uA during boiling study.
    04:28 Completed boiling test.
    04:32 Ask for 100 uA to finally begin production again.
    04:33 Changed fast FB BPM gain control back to automatic.
    04:35 Begin producton on LH2.
    05:02 Changed fast FB BPM gain control from auto to manual.

    Parity Run Summary:
    Start    Run       Events          Comments
    =======================================================================================================
    02:00    12234                     Test run for BCM/density study
    02:13    12235                     Unser switched from adc22_1 to adc23_2
    02:17    12236       1.7k          Junk, Unser switched back to adc22_1
    02:18    12237                     Test run
    02:20    12238        50k          BCM/density study ramp current 70-100
    uA, then 90,80,70,50 and 30 uA
    03:02    12239        20k          Boiling study 40 uA, BPM FB in
    automatic gain
    03:23    12240       114k          Boiling study 40, 60, 80, 100 uA, BPM
    FB in manual gain
    04:35    12241        49k          LH2 (loop 3) 100 uA BPM FB in auto gain
    05:04    12242       100k          Production LH2 (loop 3) 100 uA BPM FB
    in manual gain
    06:08    12243       100k          Production LH2 (loop 3) 100 uA BPM FB
    in manual gain
    07:07    12244       100k          Production LH2 (loop 3) 100 uA BPM FB
    in manual gain